How do you put the expression y=-2x+-5 into a table?
Answer by KMST(5328)

I like to make one column for the x values and another column for the y values, but you (or your teacher) may want each variable in one row instead.
When I make a table, it usually is a step towards making a graph, so my columns fit nicely.
Next, I pick a few values for x and calculate the corresponding y.
For example;
x = 2 means y = -2(2)+(-5) = -4 + (-5) = -9
x = 0 means y = -2(0)+(-5) = 0 + (-5) = -5
x = -3 means y = -2(-3)+(-5) = 6 + (-5) = -1
Teachers have their preferences, and may say, for example, that they want 5 values for x in the table. (It is wise to listen and follow those directions).
I place the x values , in order from smallest to greatest on the column (or row) for the x values, and I place the y values opposite the corresponding x value.

My table plus the graph that goes with it is shwon below.

I usually do my graphs on "grid paper". I buy it, or sometimes I print pages with table space and graph space. There are such thing on the internet that you can get for free. Search for "grid paper" or "graph paper."